# When Federated Learning Meets Quantum Computing: Survey and Research Opportunities

### v2

@@ 1 Introduction
Federated Learning (FL) represents a decentralized paradigm in Machine Learning (ML), introduced by Google researchers [1]. Unlike traditional centralized approaches, FL facilitates the collaborative training of a global model by multiple clients, preventing the need for local data to be transferred to a central server [2]. Clients undertake local training of models, sharing solely the model parameters with the central server for aggregation [3]. This approach ensures data privacy while facilitating learning from diverse datasets across various devices. Moreover, it capitalizes on data heterogeneity, enabling model training even in instances where data among clients are non-Independent and Identically Distributed (non-IID). Nevertheless, FL confronts technical challenges, including heterogeneity, asynchronous updates, and communication efficiency between devices and servers. As FL progresses, it is becoming an indispensable framework in privacy-preserving applications across a multitude of industries  [4, 5, 6].
@@
From a computing aspect, Quantum computing (QC) is a rapidly advancing domain that employs the principles of quantum mechanics to execute computations beyond the efficient capabilities of classical computers [7]. Central to QC is the notion of quantum bits, commonly known as qubits, which, in contrast to classical bits, possess the ability to exist in superposition(ability of a system to exist in multiple states simultaneously). This characteristic, in conjunction with entanglement [8, 9] and quantum interference [10, 11], facilitates quantum computers in executing parallel computations at an unprecedented scale, rendering them particularly adept at addressing intricate optimization problems, cryptographic challenges, and simulations of quantum physical systems [12, 13, 14, 15]. QC holds the promise to transform disciplines such as cryptography, materials science, and particularly ML. Quantum algorithms such as Shor’s algorithm [16] for factorization of large numbers and Grover’s search algorithm [17], illustrate the potential of QC to surpass classical systems in specific applications.
@@
Quantum Federated Learning (QFL) combines FL and QC to meet the growing computational demands of distributed systems and enhance privacy [18]. Traditional FL struggles with scalability and privacy, especially with large data sets and complex models. QC can address these by speeding up tasks like model optimization and parameter aggregation [19, 20]. QFL framework (as illustrated in Fig. 1) is promising for real-time data processing in edge computing and healthcare [21], offering robust, fast, and secure ML models by merging FL with efficient quantum algorithms such as Quantum Approximate Optimization Algorithms (QAOA)  [22] on Noisy Intermediate Scale Quantum (NISQ) devices [23].
@@
$\bullet$ Motivation: While several surveys on QFL exist, considerable research gaps persist, necessitating a more rigorous examination. The early work in [24] introduced a two-tier QFL model for wireless communications, but it lacks empirical validation and overlooks practical concerns such as quantum noise, scalability, and communication overhead. Later studies, such as [25, 26, 27], attempted to categorize QFL techniques but could not sufficiently establish the foundational connections between classical FL and quantum frameworks. Research efforts in [28, 29, 30, 31] have made strides in outlining the theoretical foundations of QFL, yet they do not incorporate evaluation methodologies for quantum communication constraints and real-world applicability. A recent study [29] explored QFL in the context of space-air-ground integrated networks but failed to address key hardware limitations in NISQ devices. Furthermore, [32] presented the research opportunities related to privacy aspects of QFL. Surprisingly, none of the above studies analyzes the practical implications of QFL frameworks on NISQ devices and also fails to provide a meaningful way to categorize QFL research. Thus, a more comprehensive and systematic study is needed to gain crucial insight into the potential types of QFL research along with their challenges.
@@
The remainder of the paper is organized as follows. Section 2 provides a systematic protocol for selecting relevant papers. Section 3 offers a foundation for QFL and Section 4 introduces two novel evaluation metrics along with a proposal of the novel taxonomy. Sections 567, and 8 discuss works that fall within different categories of taxonomy. We also provide insights into experimental evaluation in Section 9. Finally, Section 10 concludes the paper with open challenges and potential research directions.
@@ 2 Research Protocol
To conduct a systematic review, we follow a research protocol, inspired by the work [33], to outline the process for searching, filtering, and selecting the relevant QFL papers. This protocol offers a structured approach to uncover notable contributions that need to be explored to gain an in-depth comprehensive understanding of the field.
@@
{tblr}
cell13 = c, cell23 = c, cell24 = c, cell33 = c, cell34 = c, cell43 = c, cell44 = c, cell51 = c=2, cell54 = c, hlines, vline- = 1-4, vline1,5 = 5, Query & Query Terms Fetched
@@ 2.1 Search Strategy
Initially, we find a set of search terms to retrieve the most relevant QFL papers from well-established academic databases, including IEEE Xplore (IX), arXiv (AR), and Google Scholar (GS), as they comprehensively cover research fields pertinent to QC, FL, and QFL. Table I presents the query terms and the number of papers retrieved from each database. To ensure extensive coverage, we designed queries with varying levels of granularity while consistently including terms like quantum, federated learning, and privacy. Specific quantum terms such as ”QAOA” and ”NISQ” were also included. The search was restricted to papers published within the past 9 years (2016 to 2025) to focus on recent and relevant advancements.
@@ 3.1 Quantum Computing
QC fundamentally exploits the non-classical phenomena of superposition, interference, and entanglement to encode and process information in ways that defy classical intuition [34, 35]. As illustrated in Fig. 3, in these systems, an input quantum state $|\psi_{i}\rangle$ can exist in a superposition of basis states, enabling multiple computational pathways simultaneously, while interference modulates the probability amplitudes, enhancing desired outcomes and canceling others. Entanglement, a uniquely quantum phenomenon, correlates the states of spatially separated particles such that the overall quantum information $\phi_{i}$ becomes inseparable [36]. Table II exemplifies canonical entangled states, each illustrating different facets of quantum correlations and their critical role in quantum algorithms.
@@
Building on these foundational principles, Variational Quantum Circuits (VQC) [37] and Variational Quantum Algorithms (VQA) [38] emerge as potent paradigms that iteratively optimize parameterized unitary operators, $U(\theta)$. Here, $U(\theta)$ denotes a sequenced quantum circuit comprising layers of single-qubit rotations $R_{k}(\theta_{k})=e^{-i\theta_{k}\sigma_{k}/2}$ interleaved with entangling gates (for example, Controlled Not Gates (CNots), with the full operator taking the form
@@
where $L$ is the number of layers, $Q$ the qubit count, $R_{q}^{(\ell)}$ the rotation on qubit $q$ in layer $\ell$, and $E^{(\ell)}$ an entangling stage. They thus form smooth, differentiable mappings from a real parameter space into the unitary group $U(2^{n})$, preserving the norm of quantum states while enabling a rich family of rotations and entangling transformations [39]. Their analytic dependence on $\theta$ allows unbiased gradient evaluation via the parameter‐shift rule, making them amenable to variational optimization under NISQ limitations [40]. The circuit leverages this and transforms an input state $\lvert\psi_{i}\rangle$ into an output $\lvert\phi_{i}\rangle=U(\theta)\lvert\psi_{i}\rangle$, such that the measurement outcome $y_{i}=\langle\phi_{i}\lvert M\rvert\phi_{i}\rangle$ encodes the solution to a problem. This approach leverages quantum embeddings, which map classical data into high-dimensional Hilbert spaces (a complete inner product space that serves as the framework for representing quantum states)[41], and quantum tensors that efficiently capture the intricacies of multi-qubit interactions [42]. Quantum Neural Networks (QNNs) adopt these variational frameworks, integrating an ansatz, which is a carefully constructed circuit template, into their architecture to learn complex representations [43]. QAOA further exemplifies this strategy by embedding problem-specific constraints within its ansatz to navigate solution landscapes with heightened efficacy [44].
Yet, the promise of QC is tempered by practical challenges inherent to the NISQ era, where decoherence, gate imperfections, and qubit limitations impede flawless operation [45]. These challenges necessitate robust error mitigation strategies and innovative circuit designs to harness the true potential of quantum computation. In parallel, Fully Homomorphic Encryption (FHE) offers a classical countermeasure by enabling computations on encrypted data, thereby enhancing security in quantum communication protocols [46]. Quantum Key Distribution (QKD) further capitalizes on quantum mechanics to establish unconditionally secure channels, ensuring that cryptographic keys remain impervious to eavesdropping [47]. The interplay of these advanced concepts ranging quantum superposition and entanglement to the algorithmic sophistication of VQA and QAOA, along with the nuanced design of quantum ansatzes to the cryptographic fortifications of FHE and QKD, encapsulates the multifaceted and dynamic evolution of QC. As research continues to unravel the complexities of quantum tensors and refine variational techniques, the field strides steadily towards overcoming the limitations imposed by NISQ hardware, promising a future where quantum advantages are realized in practical, scalable systems.
@@ 3.2.1 At client
    where $\pi$ is the constant shift that leverages the periodicity of the parameterized quantum gates [34], and $L$ denotes the loss function, such as the cross-entropy loss.
@@ 3.2.2 At server
where $\eta$ is the learning rate and $w_{i}$ is the weight assigned to the $i$-th client based on its contribution. The updated global parameters $\theta^{(t+1)}$ are distributed back to the quantum clients for the next iteration. Eq. 4 resembles the classical model aggregation algorithm, known as FedAvg [1].
@@ 3.3 Advantages over FL
  • Privacy preservation: Quantum cryptographic protocols, such as quantum key exchange [47] and quantum differential privacy [48], ensure secure transmission of parameters and quantum information.
@@ 4.1.1 Qubit Utilization Efficiency (QUE)
$\mathrm{P}_{train}=\begin{cases}\mathrm{P}_{model},&\text{if specified}\\ \mathrm{L}_{model}\times\mathrm{Q}_{alloc},&\text{otherwise}\\ \end{cases}$
@@
Values near unity indicate High QUE: almost every qubit participates in every layer, signaling compact circuit depth and minimal idle hardware. Intermediate values ($0.5\leq\mathrm{QUE}\leq 0.8$) represent Moderate QUE: some qubits or layers remain unused, but the design remains reasonably dense. Lastly, values below 0.5 denote Low QUE: many qubit–layer slots sit idle, perhaps due to excessive ancilla or sparse gate placement. While no research in this survey falls under Low QUE currently, including the category would help in contrasting future QFL research on a comparable scale. For instance, the QCNN‐based QFL‐TL framework proposed in [49] uses a four‐qubit, four‐layer ansatz that achieves a QUE of 1.0, illustrating how shallow, well‐structured circuits can fully utilize qubit resources relative to gate overhead. On the other hand, the framework employed in [50], an eight‐qubit, three‐layer QCNN ansatz, yields a QUE of approximately 0.80, reflecting moderate qubit–layer occupancy relative to its entangling overhead.
This metric is particularly relevant for assessing scalability and carries particular weight on NISQ systems, where each idle qubit or under-utilized cycle magnifies decoherence risks and reduces practical throughput. QUE measurement is thus crucial, as resource-efficient designs are critical to maximizing performance.
$\bullet$ Sensitivity Analysis: Treating $\mathrm{L}_{model}$ as a tunable parameter results in significant shifts in QUE across QFL implementations. In the QCNN‐TL model [49], raising the layer count from 4 to 6 causes QUE to drop from 1.00 to 0.80. Meanwhile, the eight‐qubit embedding setup in [50] sees QUE decline from 0.80 at $\mathrm{L}_{model}=3$ to 0.60 at $\mathrm{L}_{model}=5$. Finally, in the VQA‐based scheme described in [25], doubling $\mathrm{L}_{model}$ from 4 to 8 reduces QUE from 0.75 to 0.50. These variations highlight a clear tension between circuit depth and qubit efficiency, guiding practitioners in selecting depths that balance expressivity with resource constraints.
@@ 4.1.2 Quantum Model Training Strategy (QMTS)
$\theta^{(t+1)}=\theta^{(t)}+\eta\sum_{i=1}^{K}w_{i}\,\Delta\theta_{i}^{(t)},\hskip 5.69046ptwhere\hskip 5.69046ptw_{i}=\frac{n_{i}}{\sum_{j}n_{j}}$
This rule mirrors FedAvg where each $\Delta\theta_{i}^{(t)}$ is derived from a parameter-shift quantum circuit [51]. For example, [52] implements a synchronous aggregation scheme by collecting all local variational circuit parameter updates at each round and performing a single global update. This ensures that every client’s gradient contributes equally to the model synchronization, mitigating drift and maintaining consistency across heterogeneous quantum nodes.
@@
$\theta^{(t+1)}=\theta^{(t)}+\eta\sum_{i=1}^{K}w_{i}\,\Delta\theta_{i}^{(t-\tau_{k})}$
Stale corrections can accelerate slow nodes but may introduce bias. A decay factor $\gamma^{\tau_{k}}$ can thus temper outdated gradients. This allows for staggered participation to accommodate hardware variability, which is especially important in the NISQ devices. For example, [53] employs an asynchronous aggregation, where each client pushes its encrypted quantum‐gradient message to the server. The server integrates these incoming updates on the fly, eliminating the need to wait for slower or offline participants and improving overall training.
@@
$\theta^{(t+1)}=\alpha\,\Bigl{[}\theta^{(t)}+\eta\,\sum_{i\in\mathcal{C}}w_{i}\Delta\theta_{i}\Bigr{]}+(1-\alpha)\,\Bigl{[}\theta^{(t)}+\eta\,\sum_{j\in\mathcal{Q}}w_{j}\Delta\theta_{j}\Bigr{]}$
where $0\leq\alpha\leq 1$ is used to balance the two modalities. This strategy leverages classical expressivity alongside quantum expressivity by combining quantum-classical processes to leverage the strengths of both paradigms within a unified framework. For example, [54] exemplifies a hybrid classical–quantum fusion approach. Classical clients submit standard gradient averages, while quantum clients send variational circuit parameter deltas derived from their local quantum data. The server then fuses these two modalities into a unified model update, balancing classical expressivity with quantum-enhanced feature extraction.
By formalizing these three update rules, the metric captures both the aggregation rhythm and the quantum‐classical interaction. Researchers can classify any QFL scheme by matching its update equation to either category. The importance of this metric lies mainly in its ability to reveal the adaptability of QFL frameworks in heterogeneous and dynamic environments, such as those involving uncertainty of client availability or hardware reliability. Effective training strategies ensure that federated systems maintain performance consistency while mitigating the impact of noisy operations or client dropout.

### v4

@@ 1 Introduction
Federated Learning (FL) represents a decentralized paradigm in Machine Learning (ML), introduced by Google researchers [mcmahan2017communication]. Unlike traditional centralized approaches, FL facilitates the collaborative training of a global model by multiple clients, preventing the need for local data to be transferred to a central server [FL1]. Clients undertake local training of models, sharing solely the model parameters with the central server for aggregation [FL2]. This approach ensures data privacy while facilitating learning from diverse datasets across various devices. Moreover, it capitalizes on data heterogeneity, enabling model training even in instances where data among clients are non-Independent and Identically Distributed (non-IID). Nevertheless, FL confronts technical challenges, including heterogeneity, asynchronous updates, and communication efficiency between devices and servers. As FL progresses, it is becoming an indispensable framework in privacy-preserving applications across a multitude of industries  [FLApp2, FLApp1, FLApp3].
@@
From a computing aspect, Quantum computing (QC) is a rapidly advancing domain that employs the principles of quantum mechanics to execute computations beyond the efficient capabilities of classical computers [QC1]. Central to QC is the notion of quantum bits, commonly known as qubits, which, in contrast to classical bits, possess the ability to exist in superposition(ability of a system to exist in multiple states simultaneously). This characteristic, in conjunction with entanglement [901921, 9779492] and quantum interference [philippidis1979quantum, 9716782], facilitates quantum computers in executing parallel computations at an unprecedented scale, rendering them particularly adept at addressing intricate optimization problems, cryptographic challenges, and simulations of quantum physical systems [QC3, QC4, QC5, QC6]. QC holds the promise to transform disciplines such as cryptography, materials science, and particularly ML. Quantum algorithms such as Shor’s algorithm [QC7] for factorization of large numbers and Grover’s search algorithm [QC8], illustrate the potential of QC to surpass classical systems in specific applications.
@@
Quantum Federated Learning (QFL) combines FL and QC to meet the growing computational demands of distributed systems and enhance privacy [QFL1]. Traditional FL struggles with scalability and privacy, especially with large data sets and complex models. QC can address these by speeding up tasks like model optimization and parameter aggregation [QFL2, QFL3]. QFL framework (as illustrated in Fig. 1) is promising for real-time data processing in edge computing and healthcare [QFL4], offering robust, fast, and secure ML models by merging FL with efficient quantum algorithms such as Quantum Approximate Optimization Algorithms (QAOA)  [QFL5] on Noisy Intermediate Scale Quantum (NISQ) devices [Preskill2018].
@@
$\bullet$ Motivation: While several surveys on QFL exist, considerable research gaps persist, necessitating a more rigorous examination. The early work in [narottama2020quantum] introduced a two-tier QFL model for wireless communications, but it lacks empirical validation and overlooks practical concerns such as quantum noise, scalability, and communication overhead. Later studies, such as [huang2022quantum, ren2023towards, gurung2023quantum], attempted to categorize QFL techniques but could not sufficiently establish the foundational connections between classical FL and quantum frameworks. Research efforts in [chehimi2024foundations, quy2024from, QFL-type, javeed2023quantum] have made strides in outlining the theoretical foundations of QFL, yet they do not incorporate evaluation methodologies for quantum communication constraints and real-world applicability. A recent study [quy2024from] explored QFL in the context of space-air-ground integrated networks but failed to address key hardware limitations in NISQ devices. Furthermore, [saha2024multifaceted] presented the research opportunities related to privacy aspects of QFL. Surprisingly, none of the above studies analyzes the practical implications of QFL frameworks on NISQ devices and also fails to provide a meaningful way to categorize QFL research. Thus, a more comprehensive and systematic study is needed to gain crucial insight into the potential types of QFL research along with their challenges.
@@
The remainder of the paper is organized as follows. Section 2 provides a systematic protocol for selecting relevant papers. Section 3 offers a foundation for QFL and Section 4 introduces two novel evaluation metrics along with a proposal of the novel taxonomy. Sections 567, and 8 discuss works that fall within different categories of taxonomy. We also provide insights into experimental evaluation in Section 9. Finally, Section 10 concludes the paper with recommendations for further research in QFL.
@@ 2 Research Protocol
To conduct a systematic review, we follow a research protocol, inspired by the work [gupta2020approaches], to outline the process for searching, filtering, and selecting the relevant QFL papers. This protocol offers a structured approach to uncover notable contributions that need to be explored to gain an in-depth comprehensive understanding of the field.
@@
Query Query Terms Fetched
@@ 2.1 Search Strategy
Initially, we find a set of search terms to retrieve the most relevant QFL papers from well-established academic databases, including IEEE Xplore (IX), arXiv (AR), and Google Scholar (GS), as they comprehensively cover research fields pertinent to QC, FL, and QFL. Table I presents the query terms and the number of papers retrieved from each database. To ensure extensive coverage, we designed queries with varying levels of granularity while consistently including terms like quantum, federated learning, and privacy. Specific quantum terms such as ”QAOA” and ”NISQ” were also included. The search is restricted to papers published within the past 9 years (2016 to 2025).
@@ 3.1 Quantum Computing
QC fundamentally exploits the non-classical phenomena of superposition, interference, and entanglement to encode and process information in ways that defy classical intuition [biamonte2017quantum, QFL6]. As illustrated in Fig. 3, in these systems, an input quantum state $|\psi_{i}\rangle$ can exist in a superposition of basis states, enabling multiple computational pathways simultaneously, while interference modulates the probability amplitudes, enhancing desired outcomes and canceling others. Entanglement, a uniquely quantum phenomenon, correlates the states of spatially separated particles such that the overall quantum information $\phi_{i}$ becomes inseparable [preskill2012quantum]. Table II exemplifies canonical entangled states, each illustrating different facets of quantum correlations and their critical role in quantum algorithms.
@@
Building on these foundational principles, Variational Quantum Circuits (VQC) [biswas2024sensor] and Variational Quantum Algorithms (VQA) [jin2024tetris] emerge as potent paradigms that iteratively optimize parameterized unitary operators, $U(\theta)$. Here, $U(\theta)$ denotes a sequenced quantum circuit comprising layers of single-qubit rotations $R_{k}(\theta_{k})=e^{-i\theta_{k}\sigma_{k}/2}$ interleaved with entangling gates (for example, Controlled Not Gates (CNots), with the full operator taking the form
@@
where $L$ is the number of layers, $Q$ the qubit count, $R_{q}^{(\ell)}$ the rotation on qubit $q$ in layer $\ell$, and $E^{(\ell)}$ an entangling stage. They thus form smooth, differentiable mappings from a real parameter space into the unitary group $U(2^{n})$, preserving the norm of quantum states while enabling a rich family of rotations and entangling transformations [Cerezo2021]. Their analytic dependence on $\theta$ allows unbiased gradient evaluation via the parameter‐shift rule, making them amenable to variational optimization under NISQ limitations [schuld2021effect]. The circuit leverages this and transforms an input state $\lvert\psi_{i}\rangle$ into an output $\lvert\phi_{i}\rangle=U(\theta)\lvert\psi_{i}\rangle$, such that the measurement outcome $y_{i}=\langle\phi_{i}\lvert M\rvert\phi_{i}\rangle$ encodes the solution to a problem. This approach leverages quantum embeddings, which map classical data into high-dimensional Hilbert spaces (a complete inner product space that serves as the framework for representing quantum states)[kish2003quantum], and quantum tensors that efficiently capture the intricacies of multi-qubit interactions [huggins2019towards]. Quantum Neural Networks (QNNs) adopt these variational frameworks, integrating an ansatz, which is a carefully constructed circuit template, into their architecture to learn complex representations [KAK1995259]. QAOA further exemplifies this strategy by embedding problem-specific constraints within its ansatz [mahroo2023learning].
Yet, the promise of QC is tempered by practical challenges inherent to the NISQ era, where decoherence, gate imperfections, and qubit limitations impede flawless operation [chen2024nisq]. These challenges necessitate robust error mitigation strategies and innovative circuit designs to harness the true potential of quantum computation. In parallel, Fully Homomorphic Encryption (FHE) offers a classical countermeasure by enabling computations on encrypted data, thereby enhancing security in quantum communication protocols [dutta2024federated]. Quantum Key Distribution (QKD) further capitalizes on quantum mechanics to establish unconditionally secure channels, ensuring that cryptographic keys remain impervious to eavesdropping [yang2023survey]. As research continues to unravel the complexities of quantum tensors and refine variational techniques, the field strides steadily towards overcoming the limitations imposed by NISQ hardware.
@@ 3.2.1 At client
    where $\pi$ is the constant shift that leverages the periodicity of the parameterized quantum gates [biamonte2017quantum], and $L$ denotes the loss function, such as the cross-entropy loss.
@@ 3.2.2 At server
where $\eta$ is the learning rate and $w_{i}$ is the weight assigned to the $i$-th client based on its contribution. The updated global parameters $\theta^{(t+1)}$ are distributed back to the quantum clients for the next iteration. Eq. 4 resembles the classical model aggregation algorithm, known as FedAvg [mcmahan2017communication].
@@ 3.3 Advantages over FL
  • Privacy preservation: Quantum cryptographic protocols, such as quantum key exchange [yang2023survey] and quantum differential privacy [hirche2023quantum], ensure secure transmission of parameters and quantum information.
@@
Furthermore, Quantum-enabled protocols can curb both computation and communication costs in QFL beyond raw circuit acceleration. First, federated quantum embeddings compress high-dimensional local features into amplitudes or low-qubit tensor-network states, which cuts per-round payload size and uplink time while preserving discriminative structure [Qiao2025_Comst, QFL-TL-2, ballester_quantum_2025]. Second, parameter-sparse ansätze constrain the trainable gate set and reduce forward–backward evaluations per local epoch, which shortens wall-clock time and lowers the number of parameters transmitted per round [Qiao2025_Comst]. Third, post-quantum or quantum-secure aggregation replaces heavyweight homomorphic encryption with lighter masking, which trims CPU cycles at clients and bytes on the wire [QFL-SecAggPostQ1, QFL10, Zhang2025_SAReview]. Fourth, asynchronous aggregation tolerates variable shot budgets and coherence windows, which limits idle waiting and reduces straggler amplification of both compute and message cost [Qiao2025_Comst, Chehimi2024_NSF].
@@ 4.1.1 Qubit Utilization Efficiency (QUE)
[…]
@@
Values near unity indicate High QUE: almost every qubit participates in every layer, signaling compact circuit depth and minimal idle hardware. Intermediate values ($0.5\leq\mathrm{QUE}\leq 0.8$) represent Moderate QUE: some qubits or layers remain unused, but the design remains reasonably dense. Lastly, values below 0.5 denote Low QUE: many qubit–layer slots sit idle, perhaps due to excessive ancilla or sparse gate placement. While no research in this survey falls under Low QUE currently, including the category would help in contrasting future QFL research on a comparable scale. For instance, the QCNN‐based QFL‐TL framework proposed in [QFL-TL] uses a four‐qubit, four‐layer ansatz that achieves a QUE of 1.0, illustrating how shallow, well‐structured circuits can fully utilize qubit resources relative to gate overhead. On the other hand, the framework employed in [FQE1], an eight‐qubit, three‐layer QCNN ansatz, yields a QUE of approximately 0.80, reflecting moderate qubit–layer occupancy relative to its entangling overhead.
This metric is particularly relevant for assessing scalability and carries particular weight on NISQ systems, where each idle qubit or under-utilized cycle magnifies decoherence risks and reduces practical throughput.
$\bullet$ Sensitivity Analysis: Treating $\mathrm{L}_{model}$ as a tunable parameter results in significant shifts in QUE across QFL implementations. In the QCNN‐TL model [QFL-TL], raising the layer count from 4 to 6 causes QUE to drop from 1.00 to 0.80. Meanwhile, the eight‐qubit embedding setup in [FQE1] sees QUE decline from 0.80 at $\mathrm{L}_{model}=3$ to 0.60 at $\mathrm{L}_{model}=5$. Finally, in VQA‐based scheme described in [huang2022quantum], doubling $\mathrm{L}_{model}$ from 4 to 8 reduces QUE from 0.75 to 0.50. These variations highlight a clear tension between circuit depth and qubit efficiency.
$\bullet$ Security implications of scaling qubits and gates: Beyond efficiency considerations, the implications of scaling qubit counts and circuit depth must also be examined from a security perspective. While QUE primarily captures the computational and communication overhead of scaling quantum circuits, an equally important dimension concerns the induced security risks. As the number of allocated qubits $Q_{\text{alloc}}$ and parameterized gates increases, the attack surface expands proportionally. The risk is accentuated in distributed QFL, where heterogeneous noise masks malicious parameter shifts and complicates attribution [chu2025bvqcbackdoorstylewatermarkingscheme]. Recent studies in quantum network engineering highlight that deep NISQ circuits exacerbate vulnerability to manipulation due to the limited error-correction [khalid2024quantumnetwork]. Moreover, the integration of sustainable big AI frameworks with quantum broadcast channels has underscored how expanded gate-level programmability can inadvertently widen leakage channels if not secured by anonymous or semantic communication overlays [tariq2024semantic]. For QFL deployments, the implication is clear; efficiency gains from increasing $Q_{\text{alloc}}$ must be balanced with robust monitoring of parameterized gates, anomaly detection in PQC updates, and secure aggregation mechanisms that explicitly account for gate-level tampering.
@@ 4.1.2 Quantum Model Training Strategy (QMTS)
$\theta^{(t+1)}=\theta^{(t)}+\eta\sum_{i=1}^{K}w_{i}\,\Delta\theta_{i}^{(t)},\hskip 5.69046ptwhere\hskip 5.69046ptw_{i}=\frac{n_{i}}{\sum_{j}n_{j}}.$
This rule mirrors FedAvg where each $\Delta\theta_{i}^{(t)}$ is derived from a parameter-shift quantum circuit [QFL-type]. For example, [QFL-QNGD] implements a synchronous aggregation scheme by collecting all local variational circuit parameter updates at each round and performing a single global update. This ensures that every client’s gradient contributes equally to the model synchronization and mitigating drift.
@@
$\theta^{(t+1)}=\theta^{(t)}+\eta\sum_{i=1}^{K}w_{i}\,\Delta\theta_{i}^{(t-\tau_{k})}.$
Stale corrections can accelerate slow nodes but may introduce bias. A decay factor $\gamma^{\tau_{k}}$ can thus temper outdated gradients. This allows for staggered participation to accommodate hardware variability, which is especially important in the NISQ devices. For example, [QFL-AsynchAgg] employs an asynchronous aggregation, where each client pushes its encrypted quantum‐gradient message to the server. The server integrates these incoming updates on the fly.
@@
$\theta^{(t+1)}=\alpha\,\Bigl[\theta^{(t)}+\eta\,\sum_{i\in\mathcal{C}}w_{i}\Delta\theta_{i}\Bigr]+(1-\alpha)\,\Bigl[\theta^{(t)}+\eta\,\sum_{j\in\mathcal{Q}}w_{j}\Delta\theta_{j}\Bigr],$
where $0\leq\alpha\leq 1$ is used to balance the two modalities. This strategy leverages classical expressivity alongside quantum expressivity by combining quantum-classical processes to leverage the strengths of both paradigms within a unified framework. The work [FQE2] exemplifies a hybrid classical–quantum fusion approach. Classical clients submit standard gradient averages, while quantum clients send variational circuit parameter deltas derived from their local quantum data, and then the server fuses these two modalities into a unified model update.
By formalizing these three aggregation rules, the metric captures both the aggregation rhythm and the quantum‐classical interaction. Researchers can classify any QFL scheme by matching its update equation to either category. The importance of this metric lies mainly in its ability to reveal the adaptability of QFL frameworks in heterogeneous and dynamic environments, such as those involving uncertainty of client availability or hardware reliability.